OpenStudy (anonymous):

Could someone please help me with finding the real and imaginary parts of the following complex function? Thanks!

OpenStudy (anonymous):

\[f(z)=z^{-1}\]

OpenStudy (anonymous):

(where z is assumed to be z=x+iy)

OpenStudy (dumbcow):

\[f(z) = \frac{1}{x+i y}*\frac{x-i y}{x-i y} = \frac{x-i y}{x^{2}+y^{2}}\]

OpenStudy (anonymous):

So the real part is \[x/x^2\]?

OpenStudy (anonymous):

sorry\[x/(x^2 + y^2)\]

OpenStudy (anonymous):

And imaginary is just -y?

OpenStudy (dumbcow):

yes, well -y/(x^2 +y^2)

OpenStudy (anonymous):

Oooooh I see! Thanks!

OpenStudy (dumbcow):

:)